OpenStudy (anonymous):

Please Help!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Solve these equations using the addition property. x - y = 4 x + y = 8 The lines intersect at ( __, __ )

OpenStudy (anonymous):

im going to poop in your mouth

OpenStudy (anonymous):

you need to delete your comment^

OpenStudy (anonymous):

im going to poop in your mouth, twice

OpenStudy (anonymous):

Ohhh, how I hated these. Solve for x and y to find an intersection. Pick an equation, any equation. We'll start with x-y=4. Solve for x. x=y+4. Andd, now you're stuck. So, temporarily take that as the value of y. x-4. Now plug that into your other equation for y. x+y=8. y+ (4+y)=8 Solve for y. 2y+4=8 2y=4. y=2. That is your permanent solution for y. Now you can solve for x. *The real x, not the temporary one we just used. Forget about that completely. Fun, isn't it?* We'll use the last equation again. 2+x=8. x=6. So, y= 2 and x=6. At the point where your equations equal the same on both sides, there is an intersection. 6+2=8? Yes. 6-2=4? Yes. So there's an intersection at (6,2). If that isn't right, I blame Calc II.